I. Let X be a random sample from an exponential distribution with unknown rate parameter θ and p.d.f (a) Find the probability of X> 2. (b) Find the moment generating function of X, its mean and variance. (c) Show that if X1 and X2 are two independent random variables with exponential distribution with rate parameter θ, then Y = X1 + 2 is a random variable with a gamma distribution and determine its parameters (you can use the moment generating function)
Answer #1

a) P(X > 2)

b) The moment generating function of X is

M_X(t) = E(e^{tX}) = int_{0}^{infty } e^{tx}*f_X(x)dx = int_{0}^{infty } e^{tx}* heta e^{- heta x}dx

etrdr

Rightarrow E(e^{tX})= int_{0}^{infty } heta e^{-( heta-t) x}dx

(θ-t)

(θ-t)

The mean of X is

rac{partial }{partial t }E(e^{tX})= rac{partial }{partial t}[rac{ heta}{( heta-t)}] = rac{ heta }{( heta -t)^2}

e-t)2

Rightarrow E(Xe^{tX})|_{t=0}=rac{ heta }{( heta -t)^2}|_{t=0}

Rightarrow E(X)=rac{ heta }{ heta^2}

E(X) =

Moreover,

20

Rightarrow E(X^2e^{tX})= rac{2 heta }{( heta -t)^3}

Rightarrow E(X^2e^{tX})|_{t=0}= rac{2 heta }{( heta -t)^3}|_{t=0}

20 E(X2) =

θ2

Hence, Var(X) = E(X2)-[E(X)12 = C7 92 e2 92

Var(X)2

c) Now,

The moment generating function of Y is

M_Y(t) = E(e^{t(X_1+X_2)}) = E(e^{tX_1})*E(e^{tX_2}) = rac{ heta }{( heta -t)}* rac{ heta }{( heta -t)}

(θ-t)

Since MGF uniquely identifies a distribution and the above mentioned MGF is a MGF of a gamma distribution, so we can write

Y ~Gamma(2,9)